(1 point) A 10 year $11 000 par-valued bond pays monthly coupons. If the yield rate is y 12-9% and the purchase price is $7381.84, what is the coupon rate c12? Answer:

(1 point) Two bonds, each with a face value of $13000, are redeemable at par in t-years and priced to yield y4-8%. Bond 1 of P? has a coupon rate c4-11.8% and sells for $15628.24. Bond 2 has coupon rate c-5% and sells for S R What is the value Answer: $

(1 point) An investor purchases two bonds with the following properties: Bond 1: Has a face value $1000 and is redeemable at par. Pays coupons annually at a rate of 8.3% annual and was purchased for $1097.87. Bond 2: Has a face value $1000 and is redeemable at par. Pays coupons annually at a rate of 5.8% annual and was purchased for $933.66. If both bonds mature in the same number of years and the investor yields the same rate on both bonds, find the yield rate (as a percent correct to 2 decimal places) Answer
Answer #1

Here multiple questions posted and i am solving 1st one

1)Here they are asking to find the amount received when they sold bonds , the formulae is below
Price=(coupon*(1-((1+i)^-n))/i)+(issue price*(1+i)^-n)
Coupon=maturity value*coupon rate=X
i=9%/12
n=10*12=120
issue price=11000
price=7381.84
substituing above we get coupon=36.67
Coupon rate=(36.67/11000)*12
=4%